A lift is moving up with a constant acceleration=2.5 m/s^2 .when its upward velocity is 4 a boy in the lift tosses acoin imparting it an upward velocity =5 w.r.t. himself. His fingers at the moment of toss are midway b/w the floor and the ceiling , whose total height is 2.2m.
After how much time will the coin hit the floor of the lift?
Also find the distance travelled by the coin and its displacement in the earth frame

Grade:11

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
We’ll do this with in reference frame to lift
\\a_{lift}=2.5m/s^2\vec j \\v_{lift}=4m/s\vec j \\v_{coin,rel}=5m/s\vec j \\dis_{rel}=-1.1mj \\a_{coin,rel}=-10-2.5=-12.5j \\so \\-1.1=5*t-12.5*t^2 =>t={5 \pm \sqrt{25+55} \over 25} \\$take the positive t$
for earth frame use the vel relative to earth and the time we just found out.
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty